# CCC++: Optimized Color Classified Colorization with Segment Anything Model (SAM) Empowered Object Selective Color Harmonization

### v1

@@ III-E Chromatic Diversity
Conventional color image assessment metrics often concentrate on measuring discrepancies between generated and ground truth images. However, such metrics may fall short in capturing scenarios where introducing distinctive color accents to specific regions enhances visual aesthetics. Conversely, even a modest area within an image, if strategically aligned with the color scheme of a more dominant counterpart or rendered colorless, can significantly contribute to the conventional measurement criteria. Therefore a small peak ratio (major area only) can give good performance in that criteria and minor areas are overlooked. To address the above mentioned problem, we propose a novel color image evaluation metric named Chromatic Number Ratio (CNR). The CNR quantifies the richness of color classes within the generated images compared to the ground truth images. It offers a comprehensive measure of the spectrum of colors in the generated images, enhancing our understanding of color diversity. The metric is shown in Eq. 19.
@@
where, $\mathcal{P}_{i,j}$ and $\mathcal{G}_{i,j}$ is the color class value at row $i$ and column $j$ of the generated color image $\mathcal{P}$ and Ground truth image $\mathcal{G}$. $m$ and $n$ are the image’s dimensions in the color class space. The outer summation iterates through all rows ($i$) and columns ($j$) of the image in color class space. The inner summation compares each pixel ($\mathcal{P}_{i,j}$ / $\mathcal{G}_{i,j}$) with all previous pixels in the image to check for uniqueness. $[\mathcal{P}_{i,j}=\mathcal{P}_{k,l}]$ and $[\mathcal{P}_{i,j}=\mathcal{P}_{k,l}]$ is an indicator function that returns 1 if the condition is true (if pixel values are equal) and 0 if it’s false.
@@ III-F Color Class Activation Ratio
In the context of Chromatic Number Ratio (CNR), while it effectively quantifies the class activation ratio between generated images and ground truth images, elucidating the precise count of activated color classes within the entirety of discretized color classes poses a challenge. To address the issues, we propose a novel color image evaluation metric named Color Class Activation Ratio (CCAR). The CCAR quantifies percentage of activated color classes in the colorized images . It offers a comprehensive measure of the spectrum of colors in the generated images, enhancing our understanding of color variety. The metric is shown in Eq. 20.
@@
where, $\mathcal{P}_{i,j}$ is the color class value at row $i$ and column $j$ of the generated color image $\mathcal{P}$. $m$ and $n$ are the image’s dimensions in the color class space. The outer summation iterates through all rows ($i$) and columns ($j$) of the image in color class space. The inner summation compares each pixel ($\mathcal{P}_{i,j}$) with all previous pixels in the image to check for uniqueness. $[\mathcal{P}_{i,j}=\mathcal{P}_{k,l}]$ is an indicator function that returns 1 if the condition is true (if pixel values are equal) and 0 if it is false.
@@ III-G True Activation Ratio
In the evaluation metrics CNR (Chromatic Number Ratio) and CCAR (Color Class Alignment Ratio), we effectively quantify the class activation ratio between generated and ground truth images, as well as between generated and total discretized color classes, respectively. However, determining the specific count of color classes that align precisely with ground truth images remains a challenging task. While our primary goal is to enhance the diversity and plausibility of color classes, including those belonging to minor counterparts, we also emphasize the accurate alignment of color classes, particularly focusing on preserving the presence of minor classes. For this, we propose a novel color image evaluation metric named True Activation Ratio (TAR). The TAR quantifies percentage of True activated color classes in the colorized images . It offers a comprehensive measure of the truly selected spectrum of colors in the generated images, enhancing our understanding of color variety. The metric is shown in Eq. 21.
@@
2:  Pick semantic segments ($\mathcal{S^{\text{sem}}}$) from all the segments ($\mathcal{S^{\text{all}}}$)
@@
4:     Extract $\mathcal{S}^{a}$ from $a^{*}$ using coordinates of $s$
5:     Calculate mode value ($\mathcal{M}_{a}$) of $\mathcal{S}^{a}$
6:     for each pixel ($\mathcal{P}$) in $\mathcal{S}^{a}$ do
@@
8:           Replace the value of $p$ with $\mathcal{M}_{a}$
@@
11:     Extract $\mathcal{S}^{b}$ from $b^{*}$ using coordinates of $s$
12:     Calculate mode value ($\mathcal{M}_{b}$) of $\mathcal{S}^{b}$
13:     for each pixel ($\mathcal{P}$) in $\mathcal{S}^{b}$ do
@@
15:           Replace the value of $p$ with $\mathcal{M}_{b}$
@@
19:  return  Edge harmonized a*b*
@@ IV Experiments
This study undertakes several tests to validate the suggested method in this part. Section IV-A describes the datasets used for training and testing. Section IV-B explains the specifics of implementation. Section IV-C introduces evaluation metrics. Two datasets are evaluated quantitatively in Section IV-C3 in comparison to cutting-edge techniques. Section IV-C2 presents the findings of the qualitative analysis. Ablation experiments on the network are given in Section IV-E.
     Gray            4%            6%            8%            10%            12%            14%            16%        Ground Truth
@@ IV-A1 Training
We train the proposed model using the Place365 Train dataset[32]. 1.8 million images make up the Place365 training set. 365 scene categories with approximately 5000 photos each comprise the Place365 training dataset. We took all images for training our model. Our proposed model is developed on self-supervised manner. We provide no external label of our data during train. Instead, we generate the model’s own supervisory signals or labels from the input data during training.
@@ IV-A2 Testing
Place365 Test
For testing we use the Place365 Test dataset[32]. The dataset has 328.5k images with 365 scene categories. We took all images for testing our model.
ImageNet1k val
The ImageNet1k Validation set is a subset of the ImageNet1k dataset[55]. The dataset is used widely to evaluate the performance of image classification models. It contains 50k images, spanning all 1,000 object classes of main ImageNet dataset. We took all for testing our model.
Oxford 102 Flowers test
Oxford 102 Flower[56] is a set of 102 flower categories used to classify images. The flowers are picked because they are often grow in the UK. Each class consists of a varying number of photographs, ranging from 40 to 258. Large scale, attitude, and light changes are seen in the photographs. To test our model, we used the dataset’s test set. Oxford 102 Flower Test dataset has 819 flower photos in total.
Image Celeba
The CelebFaces(CelebA)[57] dataset consists of 202,599 (2.02M) pictures, encompasses a substantial collection of facial features. The images depict individuals exhibiting various features, such as brown hair, grins, and glasses, among others. Photographs has a substantial capacity to provide extensive visual documentation, including diverse individuals, varied environmental contexts, and a multitude of positional alterations. In order to evaluate our model, a random sample of 50k photographs was chosen from the dataset.
COCO
The Common Objects in Context (COCO)[58] dataset offers an extensive collection of images annotated with object instance segmentation, object keypoints, and dense captions. This dataset encompasses tens of thousands of high-resolution images spanning multiple categories, providing a diverse and comprehensive repository for advancing research in image understanding and scene parsing. The meticulous annotations and broad scope of COCO make it an invaluable asset for researchers, enabling the exploration and development of innovative algorithms in various computer vision tasks, including object detection, segmentation, and image coloring. As a benchmark dataset, COCO plays an important role in facilitating the evaluation and comparison of state-of-the-art models and methodologies, contributing to the ongoing progress and understanding of visual recognition challenges in the field of computer vision.
@@ IV-B Implementation Set Up
The experiments were carried out on a workstation equipped with an NVIDIA GEFORCE RTX 2080 Ti graphics processing unit (GPU). The neural network was implemented using PyTorch[59] version 1.28 in Python 3.10.9. Throughout the training process, we set the batch size to 64 and utilized the Adam optimizer with a learning rate of $1\times 10^{-3}$. The momentum parameters $\beta_{1}=0.5$ and $\beta_{2}=0.999$ were employed for updating and computing the network parameters. To streamline the calculations of our loss function, each ground truth $*a*b*$ tensor was resized to a $56\times 56$ dimension. We conducted a systematic exploration of various hyperparameters and trade-off factors for our proposed model, determining their values through meticulous experimental analysis. The specific values are outlined in Tab. V.
@@ IV-B2 Testing
The batch size is set to 64 for testing purposes. To train the network for color prediction, the input picture is scaled to $224\times 224$, then fed into the network. At the conclusion of the network, the output is deconvoluted to its original size and mixed with the grayscale input $\mathbf{X}$ to produce the created color picture $\mathbf{Y}$. The average values of PSNR, SSIM, IS, and F1D are obtained for quantitative comparison using the produced images and their matching ground truth images.
@@
Gray
Without Harmonize
With Harmonize
Ground Truth
Figure 18: Results with and without color harmonization
@@ IV-C1 Evaluation Metrics
We use mean squared error (MSE), peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), learned perceptual image patch similarity (LPIPS), universal image quality index (UIQI), and frechet inception distance score (FID) to compare quantitatively our proposed model with the state-of-the-art colorization methods.
Mean Squared Error (MSE)
The MSE[60] is a widely utilized metric in the fields of image processing and machine learning. It quantifies the average squared difference between the values of an estimated or predicted image and the corresponding true or target image which is shown in Equation 22.
$MSE_{\text{RGB}}(Y,X)=\frac{1}{3N}\sum_{i=1}^{N}(Y_{i}-X_{i})^{2}$ (22)
where $N$ denotes the number of pixels, $X$ and $Y$ represent the ground truth pixel values and predicted pixel values correspondingly.
Peak Signal-to-Noise Ratio (PSNR)
The PSNR[61] is a commonly employed metric in the field of image and video processing. It serves to assess the quality of a reconstructed or compressed signal by comparing it to the original version which is shown in Equation 23. The signal-to-noise ratio (SNR) quantifies the relationship between the maximum achievable power of a signal and the power of the accompanying noise within that signal.
$PSNR=10\log_{10}\left(\frac{MAX^{2}}{MSE_{\text{RGB}}}\right)$ (23)
where $MAX=255$, the maximum pixel value in the RGB scale. The definition of MSE is shown in Equation 22.
Structural Similarity Index (SSIM)
The SSIM[61] is a commonly employed metric within the image and video processing domain which is shown in Equation 24. Its purpose is to evaluate the degree of similarity between two signals. In contrast to metrics such as MSE or PSNR, the SSIM incorporates not only the differences between individual pixels but also considers the structural information and perceptual aspects of the signals.
$SSIM(Y,X)=\frac{(2\mu_{Y}\mu_{Y}+C_{1})+(2\sigma_{YX}+C_{2})}{(\mu_{Y}^{2}+\mu_{X}^{2}+C_{1})(\sigma_{Y}^{2}+\sigma_{X}^{2}+C_{2})}$ (24)
where $Y$ is the predicted pixels and $X$ is the ground truth pixels. $\mu_{Y}$ represents the mean of $Y$ and $\mu_{X}$ represent the mean of $X$, $\sigma_{Y}^{2}$ is the variance of $Y$ , $\sigma_{X}^{2}$ is the variance of $X$ , $\sigma_{YX}$ is the covariance of $Y$ and $X$, $C_{1}$ = ${k_{1}L}^{2}$ and $C_{2}$ = ${k_{2}L}^{2}$ are constant values that are used to maintain the stability of SSIM, $L$ is the dynamic range of pixel values, $k_{1}=0.01$ and $k_{2}=0.03$.
Learned Perceptual Image Patch Similarity (LPIPS)
The LPIPS[62] is a metric utilized to quantify the perceptual similarity between images or other visual signals which is shown in Equation 25. In contrast to conventional metrics such as MSE or PSNR, the LPIPS metric incorporates higher-level perceptual characteristics and evaluates similarity based on human perception.
$d(y,x)=\sum_{l=1}^{L}\frac{1}{H_{l}W_{l}}\sum_{h,w}||(Y_{h,w}-X_{h,w})||_{2}^{2}$ (25)
where $y$ and $x$ are the predicted image and ground truth image, $d$ is the distance between $y$ and $x$, $l$ is the number of layers in the feature space, $H$ and $W$ are the height and width of feature stacks, $h$ and $w$ are the height and width of the input images and $w$ is the weight vector for converting images to features.
Universal Image Quality Index (UIQI)
The UIQI[63] is a metric utilized to evaluate the quality of an image through a comparison with a reference or ideal image. The UIQI algorithm considers both the structural and statistical characteristics of images in order to calculate a quality score which is shown in Equation 26.
$Q(Y,X)=\frac{\sigma_{YX}}{\sigma_{Y}\sigma_{X}}\times\frac{2\mu_{YX}}{\mu_{Y}^{2}+\mu_{X}^{2}}\times\frac{2\sigma_{Y}\sigma_{X}}{\sigma_{Y}^{2}+\sigma_{X}^{2}}$ (26)
where $Y$ is the predicted pixels and $X$ is the ground truth images. $\mu_{Y}$ represents the mean of $Y$ and $\mu_{X}$ represent the mean of $X$, $\sigma_{Y}^{2}$ is the variance of $Y$ , $\sigma_{X}^{2}$ is the variance of $X$ and $\sigma_{YX}$ is the covariance of $Y$ and $X$.
Fréchet Inception Distance (FID)
The FID[64] is a widely utilized metric in the domain of generative adversarial networks (GANs) for assessing the quality and diversity of generated images or samples which is shown in Equation 27. The FID is a metric used to assess the similarity between the distribution of generated samples and the distribution of real samples. The method compares the respective feature representations of generated samples and real samples.
$FID(X,Y)=|\mu_{X}-\mu_{Y}|+T_{r}(\sigma_{X}+\sigma_{Y}-2\sqrt{\sigma_{X}\sigma_{Y}})$ (27)
where $Y$ is the feature vector for the predicted image and $X$ is the feature vector for ground truth images, $\mu_{X}$ represents the mean of $X$ and $\mu_{Y}$ represent the mean of $Y$, $\sigma_{X}$ is the covariance of $X$, $\sigma_{Y}$ is the covariance of $Y$ and $T_{r}$ is the rank of the matrix.
@@ IV-C2 Qualitative Comnparison
In Figure 1, we showed a visual representation of the colorized images of our proposed model along with the gray and ground truth. From the figure, we can see that our proposed model produce very plausible and vivid color compared to the ground truth. In Figure 2, we present a set of six images to elucidate the performance disparities between regression colorization and our proposed classification-based colorization, alongside their grayscale and ground truth counterparts. A discernible distinction emerges from the visual analysis—images generated by the regression model exhibit desaturation and a pronounced grayish tone, whereas those produced by our proposed classification colorization approach showcase vibrant saturation and a diverse color palette. This visual evidence substantiates the superior color richness and diversity achieved by our proposed classification model in comparison to the desaturated outputs of the regression model.
@@
In Figure 3, six images are presented, showcasing both full classes and optimized classes colorized outputs, in conjunction with their grayscale and ground truth counterparts. The visual analysis reveals a notable distinction: full class colorization struggles to accurately capture all colors, as the increased number of classes compromises prediction accuracy. Conversely, optimized classes colorization yields more realistic and vibrant color predictions. The precision of the colorization process is enhanced when employing fewer classes, elucidating the trade-off between accuracy and complexity in the optimization of color class predictions.
In Figure 4, we present a series of six images to elucidate the outcomes of the rebalance and nobalance approaches, accompanied by their grayscale and ground truth counterparts. The visual examination reveals a substantial difference: in the nobalance approach, the accuracy of color class predictions is notably low, resulting in a desaturated visual appearance. This can be attributed to an imbalance in the number of desaturated and saturated components within an image. Conversely, the rebalance approach yields saturated and visually plausible results, demonstrating the effectiveness of addressing imbalances in color components for enhanced colorization accuracy.
@@
In Figures 5(A) and 6(B), a comprehensive comparison is presented, featuring twelve images generated by our proposed CCC++ method alongside a state-of-the-art (SOTA) classification colorization method, in conjunction with grayscale and ground truth images. The analysis from Figure 5(A) reveals that the colorization achieved by our proposed CCC++ method is characterized by more object-specific and saturated hues when compared to the SOTA classification colorization method. Notably, our method exhibits colorized images that closely resemble the ground truth, outperforming the SOTA classification colorization method. Further scrutiny in Figure 6(B) highlights the ability of our proposed method to capture a diverse range of saturated, object-specific colors, including nuanced color details in smaller objects—a distinctive strength when contrasted with the SOTA classification colorization method.
In Fig.14, we compare our proposed CCC++ method with some SOTA colorization method. From the figure we can see that, Deoldify[50] produce overall desaturated color. Iijuka[24] produce some color 3rd, 4th images but overall colorization is trends to grayish effect. Larsson[20] produce overall graish effects. CIC[30] produces some object wise color but not well saturated. Zhang[34] produces some solor in 1st, 3rd, and 6th image but overall desaturated color. Su[25] produces some color in 3rd, 4th, 6th and 8th images but desaturated in others and fully fail to generate color in 2nd and 7th images. Gain[22] produce some color in 4th, 6th and 8th images but overall desaturation. DD[65] produce some good colorization but not object wise plausible. Our proposed CCC++ method produces fully object wise plusible, saturated and vivrant colorization compared to the others and near to the ground truth.

### v2

@@ III-E Chromatic Diversity
Conventional color image assessment metrics often concentrate on measuring discrepancies between generated and ground truth images. However, such metrics may fall short in capturing scenarios where introducing distinctive color accents to specific regions enhances visual aesthetics. Conversely, even a modest area within an image, if strategically aligned with the color scheme of a more dominant counterpart or rendered colorless, can significantly contribute to the conventional measurement criteria. Therefore a small peak ratio (major area only) can give good performance in that criteria and minor areas are overlooked. To address the above mentioned problem, We introduce a new color image evaluation statistic called Chromatic Number Ratio (CNR). The CNR measures the abundance of color categories in the generated images about the ground truth photos. It thoroughly assesses the range of colors in the produced images, improving our comprehension of color variety. The metric is displayed in Equation 19.
@@
where the variables $\mathcal{P}_{i,j}$ and $\mathcal{G}_{i,j}$ represent the color class value at the specific position of row $i$ and column $j$ in the color image. The variables $\mathcal{P}$ and $\mathcal{G}$ represent the predicted image and ground truth image, respectively. The image’s dimensions in the color class space are denoted by $m$ and $n$. The outer summation iterates over all rows ($i$) and columns ($j$) of the image in the color class space. The inner summation evaluates the uniqueness of each pixel ($\mathcal{P}_{i,j}$ / $\mathcal{G}_{i,j}$) by comparing it with all preceding pixels in the image. $[\mathcal{P}_{i,j}=\mathcal{P}_{k,l}]$ and [$\mathcal{G}_{i,j}$ = $\mathcal{G}_{k,l}$] is the indicator function that evaluates to 1 when the condition is true (when pixel values are equal) and 0 when it is false.
@@ III-F Color Class Activation Ratio
Within the Chromatic Number Ratio (CNR) context, although it accurately measures the ratio of active color classes between generated images and ground truth images, determining the exact number of activated color classes within the whole set of discretized color classes is challenging. To tackle the problems, we provide a new, innovative method for assessing color images called the Color Class Activation Ratio (CCAR). The CCAR measures the proportion of activated color categories in the colorized images. It thoroughly assesses the range of colors in the produced images, improving our comprehension of color diversity. The metric is displayed in Equation 20.
@@
where, the color class value at row $i$ and column $j$ of the created color image $\mathcal{P}$ is denoted as $\mathcal{P}_{i,j}$. The picture’s dimensions in the color class space are represented by $m$ and $n$. The outer summation iterates over all rows ($i$) and columns ($j$) of the image in the color class space. The inner summation evaluates the uniqueness of each pixel ($\mathcal{P}_{i,j}$) by comparing it with all preceding pixels in the image. The equality of $\mathcal{P}_{i,j}$ and $\mathcal{P}_{k,l}$ is denoted as $[\mathcal{P}_{i,j}=\mathcal{P}_{k,l}]$, which is an indicator function that evaluates to 1 when the condition is true (when pixel values are equal) and 0 when it is false.
@@ III-G True Activation Ratio
In the evaluation metrics CNR (Chromatic Number Ratio) and CCAR (Color Class Alignment Ratio), we effectively quantify the class activation ratio between generated and ground truth images, as well as between generated and total discretized color classes, respectively. However, determining the specific count of color classes that align precisely with ground truth images remains a challenging task. While our primary goal is to enhance the diversity and plausibility of color classes, including those belonging to minor counterparts, we also emphasize the accurate alignment of color classes, particularly focusing on preserving the presence of minor classes. For this, we propose a new method for assessing color images called True Activation Ratio (TAR). The TAR measures the proportion of color classes that are accurately activated in the colorized images. It thoroughly assesses the specifically chosen range of colors in the produced images, improving our comprehension of color diversity. The metric is displayed in Equation 21.
@@
2:  Select semantic segments ($\mathcal{S}^{\text{sem}}$) from the set of all segments ($\mathcal{S}^{\text{all}}$)
@@
4:     Extract $\mathcal{S}^{a}$ from $a^{*}$ using the coordinates of $s$
5:     Calculate the mode value $\mathcal{M}_{a}$ of $\mathcal{S}^{a}$
6:     for each pixel $\mathcal{P}$ in $\mathcal{S}^{a}$ do
@@
8:           Replace the value of $\mathcal{P}$ with $\mathcal{M}_{a}$
@@
11:     Extract $\mathcal{S}^{b}$ from $b^{*}$ using the coordinates of $s$
12:     Calculate the mode value $\mathcal{M}_{b}$ of $\mathcal{S}^{b}$
13:     for each pixel $\mathcal{P}$ in $\mathcal{S}^{b}$ do
@@
15:           Replace the value of $\mathcal{P}$ with $\mathcal{M}_{b}$
@@
19:  return  Edge-harmonized $a^{*}b^{*}$
@@ IV Experiments
This study undertakes several tests to validate the suggested method in this part. Section IV-A describes the datasets used for training and testing. Section IV-B explains the specifics of implementation. Section IV-C introduces evaluation criteria. Section IV-D presents the ancient black and white photos colorization. Ablation experiments on the network are given in Section IV-E.
@@ IV-A1 Training
We train the proposed model using the Place365 Train dataset[55]. 1.8 million images make up the Place365 training set. 365 scene categories with approximately 5000 photos each comprise the Place365 training dataset. The model we propose is constructed using a self-supervised approach. During training, we do not assign any external labels to our data. Alternatively, we create the model’s supervisory signals or labels from the input data while training.
Gray     4%     6%    8%    10%    12%    14%   16%  G. Tru.
@@ IV-A2 Testing
For testing we have used 5 datasets Place365 Test[55], ImageNet1k Validation[65], Oxford 102 Flower[66], CelebFaces(CelebA)[67], and Common Objects in Context (COCO)[68].
@@ IV-B Implementation Set Up
The studies were conducted on a workstation with an NVIDIA GEFORCE RTX 2080 Ti graphics processing unit (GPU). The PyTorch version 1.28 was utilized to implement the neural network in Python 3.10.9. During the training, we selected a batch size of 64 and employed the Adam optimizer with a learning rate of $1\times 10^{-3}$. The momentum parameters $\beta_{1}=0.5$ and $\beta_{2}=0.999$ were used to update and calculate the network parameters. To simplify the computations of our loss function, we shrunk each ground truth tensor to a dimension of $56\times 56$. We investigated several hyperparameters and trade-off considerations for our suggested model, establishing their values through careful experimental analysis. The precise values are delineated in TABLE V.
@@ IV-B2 Testing
The batch size is set to 64 for testing purposes. To train the network for color prediction, the input picture is scaled to $224\times 224$, then fed into the network. At the conclusion of the network, the output is deconvoluted to its original size and mixed with the grayscale input to produce the created color picture.
@@ IV-C1 Evaluation Metrics
We conduct qualitative and quantitative comparisons between our proposed model and state-of-the-art colorization methods using various metrics, including mean squared error (MSE[70]), peak signal-to-noise ratio (PSNR[71]), structural similarity index measure (SSIM[71]), learned perceptual image patch similarity (LPIPS[72]), universal image quality index (UIQI[73]), and Fréchet inception distance score (FID[74]). These metrics thoroughly assess our model’s accuracy, perceptual quality, and overall image fidelity.
@@ IV-C2 Qualitative Comnparison
In Fig. 1, we have showed a visual representation of the colorized images of our proposed model along with the gray and ground truth. From the figure, we can see that our proposed model produce very plausible and vivid color compared to the ground truth. In Fig. 2, we have presented a set of six images to elucidate the performance disparities between regression colorization and our proposed classification-based colorization, alongside their grayscale and ground truth counterparts. A discernible distinction emerges from the visual analysis—images generated by the regression model exhibit desaturation and a pronounced grayish tone, whereas those produced by our proposed classification colorization approach showcase vibrant saturation and a diverse color palette. This visual evidence substantiates the superior color richness and diversity achieved by our proposed classification model in comparison to the desaturated outputs of the regression model.
@@
In Fig. 3, six images have been presented, showcasing both full classes and optimized classes colorized outputs, in conjunction with their grayscale and ground truth counterparts. The visual analysis reveals a notable distinction: full class colorization struggles to accurately capture all colors, as the increased number of classes compromises prediction accuracy. Conversely, optimized classes colorization yields more realistic and vibrant color predictions. The precision of the colorization process is enhanced when employing fewer classes, elucidating the trade-off between accuracy and complexity in the optimization of color class predictions.
Gray
W/o Har.
W. Har.
G. Truth
Figure 18: Results with and without color harmonization
In Fig. 4, we present a series of six images to elucidate the outcomes of the rebalance and nobalance approaches, accompanied by their grayscale and ground truth counterparts. The visual examination reveals a substantial difference: in the nobalance approach, the accuracy of color class predictions is notably low, resulting in a desaturated visual appearance. This can be attributed to an imbalance in the number of desaturated and saturated components within an image. Conversely, the rebalance approach yields saturated and visually plausible results, demonstrating the effectiveness of addressing imbalances in color components for enhanced colorization accuracy.
@@ G. Truth
In Fig.’s 5(A) and 6(B), a comprehensive comparison is presented, featuring twelve images generated by our proposed CCC++ method alongside a state-of-the-art (SOTA) classification colorization method, in conjunction with grayscale and ground truth images. The analysis from Fig. 5(A) reveals that the colorization achieved by our proposed CCC++ method is characterized by more object-specific and saturated hues when compared to the SOTA classification colorization method. Notably, our method exhibits colorized images that closely resemble the ground truth, outperforming the SOTA classification colorization method. Further scrutiny in Fig. 6(B) highlights the ability of our proposed method to capture a diverse range of saturated, object-specific colors, including nuanced color details in smaller objects—a distinctive strength when contrasted with the SOTA classification colorization method.
In Fig.14, we compare our proposed CCC++ method with some SOTA colorization method. From the figure we can see that, Deoldify[60] produce overall desaturated color. Iijuka[28] produce some color 3rd, 4th images but overall colorization is trends to grayish effect. Larsson[24] produce overall graish effects. CIC[33] produces some object wise color but not well saturated. Zhang[13] produces some color in 1st, 3rd, and 6th image but overall desaturated color. Su[29] produces some color in 3rd, 4th, 6th and 8th images but desaturated in others and fully fail to generate color in 2nd and 7th images. Gain[26] produce some color in 4th, 6th and 8th images but overall desaturation. DD[41] produce some good colorization but not object wise plausible. Our proposed CCC++ method produces fully object wise plausible, saturated and vibrant colorization compared to the others and near to the ground truth.